OpenStudy (anonymous):

What is the solution to the equation? 5(8 + 7) = 5 • x + 5 • 7 A. x = 1 B. x = 5 C. x = 7 D. x = 8

OpenStudy (anonymous):

X=8 5(8+7)=5*x+5*7 bracket common terms 5(8+7)=5(x+7) X must = 8